# CaReAQA: A Cardiac and Respiratory Audio Question Answering Model for Open-Ended Diagnostic Reasoning

### v1

@@ 2.1 General-Purpose Audio-Language Models
Recent advances in audio-language models have significantly enhanced the integration of auditory perception and reasoning, enabling LLMs to handle diverse audio-related tasks. For instance, the LTU (Listen, Think, Understand) model (Gong et al., 2024) excels in both closed and open-ended question-answering tasks.
Instruction-based frameworks like Pengi (Deshmukh et al., 2023) improve task versatility by framing diverse tasks as instruction-response problems, while models like GAMA (et al., 2024c) and AudioGPT (Huang et al., 2023) leverage pre-trained audio models for reasoning over complex auditory inputs. AudioGPT, in particular, excels in multi-turn dialogues for general audio tasks but relies on predefined interfaces like ChatGPT and external systems (e.g., ASR or task-specific models) for audio processing.
@@ 2.2 Domain-Specific Audio and Multimodal Applications
General multimodal large language models, such as Gemini (Gemini, 2024), QWen-Audio (et al., 2023) and GPT-4o (OpenAI, 2024), extend reasoning capabilities to include vision, language, and auditory modalities. However, challenges remain in addressing domain-specific auditory tasks that require specialized expertise. Our approach emphasizes the use of medical audio inputs to provide open-ended answers, thereby supporting diagnostic decision-making in healthcare scenarios.
@@ 3.2 CaReSound Benchmark Dataset
  • CirCor (Oliveira, 2022): Pediatric heart sound recordings from 1,568 subjects, with detailed annotations of murmurs and heart sound segmentation.
  • SPRSound (Zhang, 2024): A pediatric respiratory sound database of annotated respiratory recordings from 292 participants. It includes detailed annotations at both record and event levels for classifications of respiratory sounds.
@@ 3.3.1 Open-Ended Question Answering
The evaluation assesses the model’s ability to generate accurate and clinically relevant diagnostic answers. We use BERTScore (Zhang* et al., 2020), a metric that quantifies semantic similarity between generated responses and reference answers. Implementation of BERTScore leverages Hugging Face’s “evaluate” library with the pre-trained “bert-base-uncased model”. To measure lexical similarity while accounting for stemming, synonyms, and word order, we include METEOR (Lavie and Agarwal, 2007) as an additional metric. In some cases, an incorrect diagnosis may have a high similarity with the correct one, due to subtle lexical differences, negations, etc. To circumvent this issue, we compute an Accuracy score by comparing the model’s predictions against the ground truth using GPT-4o (OpenAI, 2024). We prompt the GPT-4o model via API to compare whether the prediction and ground truth are the same, providing a binary judgment “Yes” or “No” for each pair, indicating whether the generated answer and the reference are semantically identical. Details of the evaluation prompt used with GPT-4o can be found in Appendix B. We aggregate these binary judgments to compute the model’s overall accuracy.
@@ 3.3.2 Generalization to Unseen Data
BMD-HS (et al., 2024b), featuring heart sound recordings from 59 patients, annotated into six categories, including valvular heart diseases such as aortic stenosis and mitral regurgitation, as well as healthy (normal) cases (6 classes).
@@ 4 Experiments
Qwen2-Audio (et al., 2024a) $47.2$ $23.4$ $17.1$ $48.2$ $29.8$ $18.8$ $34.6$ $0.3$ $7.9$ $48.3$ $25.4$ $32.7$ $49.5$ $26.8$ $19.0$ $53.5$ $30.2$ $17.0$
Gama (et al., 2024c) $59.4$ $27.5$ $17.9$ $62.2$ $30.6$ $33.6$ $39.1$ $0.3$ $5.9$ $60.3$ $29.5$ $16.6$ $64.6$ $34.2$ $26.3$ $56.8$ $28.4$ $18.2$
@@ 4.1 Training Recipe
During training, audio inputs are segmented into 5-second clips on the fly. For each batch, a random 5-second segment is extracted from each audio clip to account for variability in recording durations, ensuring consistent input lengths across batches. To enhance dataset diversity, we apply data augmentation using the AugLy library (Papakipos and Bitton, 2022). Augmentations are randomly and selectively applied to individual audio clips within each batch, introducing variability while preserving diagnostic features. Specifically, we apply one of four transformations with equal probability: a 5dB volume increase, amplitude normalization, a low-pass filter (cutoff at 300Hz), or a high-pass filter (cutoff at 3000Hz).
@@ 4.2 Open-Ended Audio Task Experiments
Qwen2-Audio (et al., 2024a) 40.2 52.5 51.0 43.1 46.7
Gama (et al., 2024c) 33.8 30.3 29.8 27.5 30.3
@@ Appendix E Comparison of Instruction-Tuned Language Models
We compared multiple large language models—namely gemma-2b-it (Gemma Team, 2024), Qwen2-1.5B-Instruct(et al., 2024a), and SmolLM2-1.7B-Instruct(Allal et al., 2024), Llama-3.2-3B-Instruct(Touvron et al., 2023)—to the base Llama-3.2-3B under identical fine-tuning conditions tailored to audio-QA. As shown in Table 12, while gemma-2b-it demonstrated occasional benefits, the base model consistently matched or exceeded the performance of all instruction-tuned options once domain-specific training was applied. These results indicate that directly refining the core model can be more effective than relying on broad instruction-tuning strategies for specialized tasks. Consequently, we adopted the base Llama-3.2-3B for CaReAQA.

### v2

@@ 2.1 General-Purpose Audio-Language Models
Recent advances in audio-language models have significantly enhanced the integration of auditory perception and reasoning, enabling LLMs to handle diverse audio-related tasks. For instance, the LTU (Listen, Think, Understand) model (Gong et al., 2023) excels in both closed and open-ended question-answering tasks.
Instruction-based frameworks like Pengi (Deshmukh et al., 2023) improve task versatility by framing diverse tasks as instruction-response problems, while models like GAMA (Ghosh et al., 2024) and AudioGPT (Huang et al., 2024) leverage pre-trained audio models for reasoning over complex auditory inputs. AudioGPT, in particular, excels in multi-turn dialogues for general audio tasks but relies on predefined interfaces like ChatGPT and external systems (e.g., ASR or task-specific models) for audio processing.
@@ 2.2 Domain-Specific Audio and Multimodal Applications
General multimodal large language models, such as Gemini (Gemini, 2024), QWen-Audio (Chu et al., 2023) and GPT-4o (OpenAI, 2024), extend reasoning capabilities to include vision, language, and auditory modalities. However, challenges remain in addressing domain-specific auditory tasks that require specialized expertise. Our approach emphasizes the use of medical audio inputs to provide open-ended answers, thereby supporting diagnostic decision-making in healthcare scenarios.
@@ 3.2 CaReSound Benchmark Dataset
  • CirCor (Oliveira et al., 2021): Pediatric heart sound recordings from 1,568 subjects, with detailed annotations of murmurs and heart sound segmentation.
  • SPRSound (Zhang et al., 2022): A pediatric respiratory sound database of annotated respiratory recordings from 292 participants. It includes detailed annotations at both record and event levels for classifications of respiratory sounds.
@@ 3.3.1 Open-Ended Question Answering
The evaluation assesses the model’s ability to generate accurate and clinically relevant diagnostic answers. We use BERTScore (Zhang et al., 2019), a metric that quantifies semantic similarity between generated responses and reference answers. Implementation of BERTScore leverages Hugging Face’s “evaluate” library with the pre-trained “bert-base-uncased model”. To measure lexical similarity while accounting for stemming, synonyms, and word order, we include METEOR (Banerjee and Lavie, 2005) as an additional metric. In some cases, an incorrect diagnosis may have a high similarity with the correct one, due to subtle lexical differences, negations, etc. To circumvent this issue, we compute an Accuracy score by comparing the model’s predictions against the ground truth using GPT-4o (OpenAI, 2024). We prompt the GPT-4o model via API to compare whether the prediction and ground truth are the same, providing a binary judgment “Yes” or “No” for each pair, indicating whether the generated answer and the reference are semantically identical. Details of the evaluation prompt used with GPT-4o can be found in Appendix B. We aggregate these binary judgments to compute the model’s overall accuracy.
@@ 3.3.2 Generalization to Unseen Data
BMD-HS (Ali et al., 2024), featuring heart sound recordings from 59 patients, annotated into six categories, including valvular heart diseases such as aortic stenosis and mitral regurgitation, as well as healthy (normal) cases (6 classes).
@@ 4 Experiments
Qwen2-Audio (Chu et al., 2023) $47.2$ $23.4$ $17.1$ $48.2$ $29.8$ $18.8$ $34.6$ $0.3$ $7.9$ $48.3$ $25.4$ $32.7$ $49.5$ $26.8$ $19.0$ $53.5$ $30.2$ $17.0$
Gama (Ghosh et al., 2024) $59.4$ $27.5$ $17.9$ $62.2$ $30.6$ $33.6$ $39.1$ $0.3$ $5.9$ $60.3$ $29.5$ $16.6$ $64.6$ $34.2$ $26.3$ $56.8$ $28.4$ $18.2$
@@ 4.1 Training Recipe
During training, audio inputs are segmented into 5-second clips on the fly. For each batch, a random 5-second segment is extracted from each audio clip to account for variability in recording durations, ensuring consistent input lengths across batches. To enhance dataset diversity, we apply data augmentation using the AugLy library (DBLP:journals/corr/abs-2201-06494). Augmentations are randomly and selectively applied to individual audio clips within each batch, introducing variability while preserving diagnostic features. Specifically, we apply one of four transformations with equal probability: a 5dB volume increase, amplitude normalization, a low-pass filter (cutoff at 300Hz), or a high-pass filter (cutoff at 3000Hz).
@@ 4.2 Open-Ended Audio Task Experiments
Qwen2-Audio (Chu et al., 2023) 40.2 52.5 51.0 43.1 46.7
Gama (Ghosh et al., 2024) 33.8 30.3 29.8 27.5 30.3
@@ Appendix E Comparison of Instruction-Tuned Language Models
We compared multiple large language models—namely gemma-2b-it (Gemma Team, 2024), Qwen2-1.5B-Instruct(Chu et al., 2023), and SmolLM2-1.7B-Instruct(Allal et al., 2024), Llama-3.2-3B-Instruct(Touvron et al., 2023)—to the base Llama-3.2-3B under identical fine-tuning conditions tailored to audio-QA. As shown in Table 12, while gemma-2b-it demonstrated occasional benefits, the base model consistently matched or exceeded the performance of all instruction-tuned options once domain-specific training was applied. These results indicate that directly refining the core model can be more effective than relying on broad instruction-tuning strategies for specialized tasks. Consequently, we adopted the base Llama-3.2-3B for CaReAQA.